# Experimental observation of gapped shear waves and liquid-like to gas-like dynamical crossover in active granular matter

### v1

@@ Introduction
In summary, the analysis of the VDOS and the atomic displacements reveal a continuous transition from a gas-like behavior typical of dilute liquids to a collective and viscoelastic motion characteristic of dense liquids by increasing the packing fraction. This perfectly aligns with the structural and dynamical transition between a gas-like liquid and a rigid liquid phase discussed in the previous section and displayed in Fig.2. Notably, the study of collective modes performed in Fig.5 confirms that $\phi$ plays the role of the inverse temperature of classical thermal liquids (see Fig.3 and related discussion above). Indeed, the behavior of the VDOS shown in panels (a)-(d) in Fig.4 is perfectly compatible with that found in liquids upon decreasing $T$ (see, for example, Fig.6 in [65]).
@@
Figure 5: Emergence of wave-vector gapped shear waves in liquid granular matter. The longitudinal (blue) and transverse (red) dynamical structure factors $S_{L,T}(\omega,k)$ for packing fraction $\phi=0.618$, $\phi=0.773$, $\phi=0.811$ and $\phi=0.822$ from left to right. $d_{s}$ is the diameter of small particles. In panels (c) and (d), the solid black lines show the fitting to the linear dispersion relation $\mathrm{Re}(\omega)=v_{L}k$. The fitted longitudinal speed of sound is $v_{L}=101.2$ and $v_{L}=116.9$ respectively in unit of $(d_{s}/s)$. The colored bullets indicate the peak position obtained by fitting $S_{L}(\omega,k)$ at fixed values of the wave vector $k$. In panels (g) and (h), the numerical data are fitted with the $k$-gap formula $\mathrm{Re}(\omega)=\sqrt{v_{T}^{2}(k^{2}-k_{g}^{2})}$ where $k_{g}$ is the size of the $k$-gap. We obtain a $k$-gap of $0.093$ for $\phi=0.811$ (g) and $0.068$ for $\phi=0.822$ (h) in units of $d_{s}^{-1}$. The fitted transverse speed of sound is $v_{T}=48.5$ and $v_{T}=71.7$ respectively in unit of $(d_{s}/s)$. In all panels, the error bar indicates the linewidth of the corresponding peak in the dynamic structure factor.
As we move to higher packing fraction data, $\phi=0.811,0.822$ shown in panels (c)-(g) and (d)-(h), the fingerprints of liquid-like collective modes become more pronounced. By fitting the dynamical structure factor, we have extracted the frequency $\Omega(k)$ and the line-width $\Gamma(k)$ of the lowest collective modes in both the longitudinal and transverse sectors. Details regarding the fitting process can be found in the Methods section, and additional figures can be found in the Supplementary Information section. In the longitudinal sector, as shown in panels (c) and (d), we observe a mode that disperses linearly at large wavelengths, with $\Omega_{L}(k)=v_{L}k$, where $v_{L}$ is of the order $10^{2}$ in units of $d_{s}$/s. The speed increases slowly as the packing fraction increases, confirming that $\phi$ plays the role of inverse temperature in classical thermal liquids. At $k\approx 0.1\,d_{s}^{-1}$, the dispersion bends down towards a constant pseudo-Van-hove plateau. Additionally, the longitudinal sound mode has a linewidth that increases with the wave vector $k$ and becomes overdamped for small wavelengths.
@@
Interestingly, the size of the $k$-gap can vary from Å in classical liquids to mm in dusty plasmas. However, its value is always an $\mathcal{O}(1)$ number in units of the Wigner-Seitz radius. The table and figures in [68] provide more information on this. Our experimental observations, as shown in Fig.5, are consistent with this trend. This suggests that the $k$-gap size may be universal in all liquids, regardless of their microscopic nature and interactions.
Finally, the lifetime of local connectivity, or $\tau_{LC}$, in classical liquids, is closely related to the Maxwell relaxation time, or $\tau$, as per a research study [69]. For a value of $\phi=0.822$, our analysis yields a value of $\tau_{LC}$ as 0.3 seconds, which is in proximity to the Maxwell relaxation time of 0.11 seconds, calculated from the size of $k$-gap (as shown in panel (d) of Fig.5). This similarity in values confirms the validity of our data analysis method for obtaining the dynamical structure factor. Also, it suggests that classical liquids in the supercritical state and granular fluids share similarities at the smallest particle level.
@@ Conclusion
Our study demonstrates that, despite significant differences in the size of composed particles and the absence of a classical thermodynamic description, granular matter on the Brownian vibrator exhibits various similarities with classical liquids in the supercritical region both in the collective and particle-level motions. These alike aspects include pair correlation functions, velocity autocorrelation, mean square displacement on particle levels, vibrational density of states, and dispersion relation of collective excitations. With an increase in the packing fraction $\phi$, distinct phases from gas-like to condensed states (liquid or solid) have been observed on both particle and collective motion levels, suggesting the role of $\phi$ as an effective inverse temperature variable, that is corroborated by the anti-correlation between the average kinetic energy and the packing fraction. Furthermore, our work experimentally revealed the emergence of the $k$-gap, as predicted by viscoelasticity theory, in granular fluids. This provides a direct link between classical liquid and active granular matter and experimentally confirms the phenomenon of $k$-gap and its theoretical premises.
Methods
@@
In this setup, the frequency $f$ is set at 100 Hz, and the maximum acceleration $a$ is 2.5 times the gravitational acceleration $g=9.8m/s^{2}$. The vibration amplitude $A$, defined as $a/(2\pi f)^{2}$, is 0.062 mm, significantly smaller than the particle’s vertical dimension (6 mm). Therefore, we can neglect the vertical displacement of a particle, treating the system as quasi-two-dimensional [49].
@@
The particles are placed on top of an aluminum plate (60 cm × 60 cm) and confined within a flower-shaped boundary designed to suppress creep particle motions along the boundary [49, 70]. We initiate the experiment with all particles randomly and uniformly placed on the base plate. After applying vibration for two hours, we achieve an initial state of particle configuration. Subsequently, the particle layer undergoes continuous vertical vibration, while a Basler CCD camera (acA2040-180kc) records particle motion at 40 frames/s for at least an hour.
@@
The following fitting functions have been used to analyze the experimentally obtained dynamical structure factor (see additional figures in the Supplementary Information). The transverse and longitudinal components of the dynamical structure factor are respectively fitted using
@@
The first terms are simply a damped harmonic oscillator with energy $\Omega_{L,T}(k)$ and linewidth $\Gamma_{L,T}(k)$. The second term is a quasi-elastic contribution modelled with a Lorentzian function. The dispersion relations $\Omega_{T,L}(k)$ are shown in Fig.5 as colored bullets and the corresponding linewidths $\Gamma_{T,L}$ by the related errors bars. $S_{\text{loc}}(\omega,k)$ is a term representing possible quasi-localized low-energy modes [72], and it is not relevant for the present discussion. More details on the fitting procedure and the raw data for the dynamical structure factors can be found in the Supplementary Information.
Supplementary information
Debye normalized density of states
@@
$g_{\text{Debye}}(\omega)=\frac{L^{2}}{2\pi\bar{v}^{2}}\,\omega\qquad\text{with}\qquad\frac{2}{\bar{v}^{2}}=\frac{1}{v_{T}^{2}}+\frac{1}{v_{L}^{2}}.$ (16)
The reduced VDOS $g(\omega)/\omega$ is presented for the two highest values of the packing fraction in Fig.6.
For the highest packing fraction, $\phi=0.822$, the longitudinal component of the reduced VDOS approaches a constant plateau at low frequency, implying the validity of the Debye model. On the contrary, the transverse component follows Debye’s law only between approximately $10$ Hz to $5$ Hz and then strongly deviates from it below such a frequency. This manifests a residual quasi-elastic diffusive contribution and a finite value of $g_{T}(\omega)$ at zero frequency. By extracting the speeds of longitudinal and transverse sound from the dispersion relation of the collective modes, we have verified that the ratio of the two plateau values in the reduced VDOS (dashed horizontal lines in Fig.6) is compatible with the theoretical prediction of Debye theory, namely $v_{T}^{2}/v_{L}^{2}$.
Figure 6: The reduced VDOS for packing fraction $\phi=0.811$ (a) and $\phi=0.822$ (b). The blue lines indicate the longitudinal component, while the red and the gray indicate respectively the transverse and total ones. The dashed lines show the theoretical predictions from the Debye model as explained in the text.
The situation becomes more complex by decreasing the packing fraction, as shown in the left panel of Fig.6 for $\phi=0.811$. There, Debye’s model no longer provides a good description of the experimental data. We can only observe a small plateau in the longitudinal component of the reduced VDOS around $5$ Hz, which is then modified at lower frequencies. In general, by increasing the system’s fluidity or equivalently decreasing its packing fraction, we no longer expect the Debye model to be a good representation of the collective dynamics.
Dynamical structure factor
We provide more detail about the analysis of the experimental dynamical structure factor. The longitudinal dynamical structure factor is fitted using a single damped harmonic oscillator,
$S_{L}(\omega,k)\propto\dfrac{\omega^{2}}{(\omega^{2}-\Omega_{L}^{2})^{2}+\omega^{2}\Gamma_{L}^{2}},$ (17)
@@
$S_{T}(\omega,k)\propto\dfrac{\omega^{2}}{(\omega^{2}-\Omega_{T}^{2})^{2}+\omega^{2}\Gamma_{T}^{2}}+\dfrac{1}{\pi}\dfrac{\gamma}{\omega^{2}+\gamma^{2}}+S_{loc}(\omega,k),$ (18)
@@
$S_{loc}(\omega,k)=\dfrac{1}{\omega\sigma\sqrt{2\pi}}\mathrm{exp}\left(-\dfrac{(\mathrm{ln}\omega-\mu)^{2}}{2\sigma^{2}}\right).$ (19)
@@
Displacement field power spectrum
@@
We have performed the Fourier transform on the displacement field to observe its evolution with time. Fig. 10 presents the results for different packing fractions. For low packing fractions (a) and (b), all Fourier transform components concentrate strongly and stably at the low wave vector region after a long observation time (800s). This indicates that the particles do not return to their initial position. However, the features differ for high packing fractions, as shown in panels (c) and (d). Firstly, the Fourier components are unstable, especially the zig-zag structures at $k\approx 0$. This suggests that the particles may still return to their initial position. Secondly, the Fourier transform components concentrate at a wave vector smaller than $k$-gap. This correlation can be interpreted as follows: the modes with $k>k_{g}$ are solid-like, meaning that the particles only vibrate around their equilibrium position, and therefore, these modes do not contribute to the displacement at late time. On the other hand, the modes with $k<k_{g}$ are gas- or liquid-like, meaning that the particles can diffuse away from their position and never return. Therefore, only these modes contribute to the long-time displacement.
Lifetime of local atomic connectivity
@@
We calculate the lifetime of local atomic connectivity, $\tau_{LC}$, which represents the average time it takes for atoms to lose or gain one nearest neighbor, changing the local atomic connectivity through a local configurational excitation. The Voronoi cell of the particle at time $t=t_{0}$ defines its nearest neighbors. The schematic diagram of Fig.11 indicates that any two particles that share an edge of a Voronoi cell are considered neighbors, as explained in detail in [73]. By doing so, the atomic bonds of a particle with its neighbors can be defined. After that, we can detect the formation or breaking of atomic bonds for $t>t_{0}$ by identifying the moments when the number of neighbors changes by a unity. For each atom, $t_{LC}$ is the time to change one neighbor such that $N_{c}(t_{0}+t_{LC})=N_{c}(t_{0})\pm 1$. The lifetime of local atomic connectivity is then defined as the average of $t_{LC}$, which is denoted as $\tau_{LC}=\langle t_{LC}\rangle$. The distribution of $t_{LC}$ for different packing fraction is shown in Fig.12. $\tau_{LC}$ values for $\phi$ values 0.618, 0.773, 0.811, and 0.822 are 0.308, 0.609, 0.507, and 0.300s respectively.
@@
Figure 7: Transverse (red,a) and longitudinal (blue,b) dynamical structure factor at different wave vector for $\phi=0.811$. The thinner lines in transverse panel show the three terms in the fitting function.
@@
Figure 8: Transverse (red,a) and longitudinal (blue,b) dynamical structure factor at different wave vector for $\phi=0.822$. The thinner lines in transverse panel show the three terms in the fitting function.
@@
Figure 9: The fitting parameters $\Omega$ (circle) and $\Gamma$ (square) for longitudinal (blue) and transverse (red) dynamical structure factor and for packing fraction $\phi=0.811$ (a) and $\phi=0.822$ (b).
@@
Figure 10: Fourier transform on displacement field, $\boldsymbol{D}(\boldsymbol{k})=\mathcal{F}[\boldsymbol{D}(\boldsymbol{x})]$, for packing fraction $\phi=0.618$, $\phi=0.773$, $\phi=0.811$ and $\phi=0.822$ from (a) to (d) respectively. Since our system have rotational symmetry, only $x$ component $D_{x}(\boldsymbol{x})$ are plot. The darker color indicate stronger Fourier transform signal. The red lines in panels (c) and (d) show the position of $k$-gap.
Figure 11: Schematic diagram for an atom to change one neighbor.
Figure 12: Distribution of $t_{LC}$ for packing fraction $\phi=0.618$, $\phi=0.773$, $\phi=0.811$ and $\phi=0.822$ from (a) to (d). $\tau_{LC}=0.308,0.609,0.507,0.300s$ for each packing fraction respectively.

### v2

@@ Introduction
In summary, the analysis of the VDOS and the particle displacements reveal a continuous transition from a gas-like behavior typical of dilute liquids to a collective and viscoelastic motion characteristic of dense liquids by increasing the packing fraction. This perfectly aligns with the structural and dynamical transition between a gas-like liquid and a rigid liquid phase discussed in the previous section and displayed in Fig.2. The study of collective modes performed in Fig.5 confirms that $\phi$ plays the role of the inverse temperature of classical thermal liquids (see Fig.3 and related discussion above). Indeed, the behavior of the VDOS shown in panels (a)-(d) in Fig.4 is perfectly compatible with that found in liquids upon decreasing $T$ (see, for example, Fig.6 in [72]).
Figure 5: Emergence of wave-vector gapped shear waves in liquid granular matter. The longitudinal (blue) and transverse (red) dynamical structure factors $S_{L,T}(\omega,k)$ for packing fraction $\phi=0.618$, $\phi=0.773$, $\phi=0.811$ and $\phi=0.822$ from (a) to (d) on top and from (e) to (h) on bottom. $d_{s}$ is the diameter of small particles. In panels (c) and (d), the solid black lines show the fitting to the linear dispersion relation $\mathrm{Re}(\omega)=v_{L}k$. The fitted longitudinal speed of sound is $v_{L}=101.2$ and $v_{L}=116.9$ respectively in unit of $(d_{s}/s)$. The colored bullets indicate the peak position obtained by fitting $S_{L}(\omega,k)$ at fixed values of the wave vector $k$. In panels (g) and (h), the numerical data are fitted with the $k$-gap formula $\mathrm{Re}(\omega)=\sqrt{v_{T}^{2}(k^{2}-k_{g}^{2})}$ where $k_{g}$ is the size of the $k$-gap. We obtain a $k$-gap of $0.093$ for $\phi=0.811$ (g) and $0.068$ for $\phi=0.822$ (h) in units of $d_{s}^{-1}$. The fitted transverse speed of sound is $v_{T}=48.5$ and $v_{T}=71.7$ respectively in unit of $(d_{s}/s)$. In all panels, the error bar indicates the linewidth of the corresponding peak in the dynamic structure factor. The dynamical structure factor is dimensionless and normalized by its maximum value, hence, the color bar goes from 00 to $1$ in all panels.
@@
As we move to higher packing fraction data, $\phi=0.811,0.822$ shown in panels (c)-(g) and (d)-(h), the fingerprints of liquid-like collective modes become more pronounced. By fitting the dynamical structure factor, we have extracted the frequency $\Omega(k)$ and the line-width $\Gamma(k)$ of the lowest collective modes in both the longitudinal and transverse sectors. Details regarding the fitting process can be found in the Methods section, and additional figures can be found in the Supplementary note 3. In the longitudinal sector, as shown in panels (c) and (d), we observe a mode that disperses linearly at large wavelengths, with $\Omega_{L}(k)=v_{L}k$, where $v_{L}$ is of the order $10^{2}$ in units of $d_{s}$/s. The speed increases slowly as the packing fraction increases, confirming that $\phi$ plays the role of inverse temperature in classical thermal liquids. At $k\approx 0.1\,d_{s}^{-1}$, the dispersion bends down towards a constant pseudo-Van-hove plateau. Additionally, the longitudinal sound mode has a linewidth that increases with the wave vector $k$ and becomes overdamped for small wavelengths.
@@
$\phi$ $v_{L}/v_{\mathrm{th}}$ $v_{T}/v_{\mathrm{th}}$
$0.811$ $86.0$ $41.2$
$0.822$ $121.2$ $74.3$
Table 1: Speed of sound. The sound velocities obtained from the experimental data in units of the thermal velocity $v_{\mathrm{th}}$ derived from $Z(0)$ of all particles (see Fig.3).
In simple fluids with steep interactions, it has been found that $v_{L}\approx 10v_{\mathrm{th}}$ and $v_{T}\approx 5v_{\mathrm{th}}$ near freezing [75], where $v_{\mathrm{th}}$ is the thermal velocity of the system. Using our experimental data, we have estimated the ratio between the sound velocities and the thermal velocity computed from $Z(0)$ (see Table 1). We found that both ratios are one order of magnitude larger than in simple fluids. We speculate that this difference is a direct consequence of the activity of the system and of the strong deviations from local thermal equilibrium.
The lifetime of local connectivity, or $\tau_{LC}$, in classical liquids, is closely related to the Maxwell relaxation time, or $\tau$, as per a research study [68]. For a value of $\phi=0.822$, our analysis yields a value of $\tau_{LC}$ as 0.3 seconds, which is in proximity to the Maxwell relaxation time of 0.11 seconds, calculated from the size of $k$-gap (as shown in panel (d) of Fig.5). This similarity in values confirms the validity of our data analysis method for obtaining the dynamical structure factor. Also, it suggests that classical liquids in the supercritical state and granular fluids share similarities at the smallest particle level.
To understand how various parameters in the experimental setup is also necessary, in addition to the packing fraction $\phi$, influence the $k$-gap behavior. We conducted additional experiments to investigate this by increasing the driving frequency from 100 Hz to 130 Hz while maintaining the highest packing fraction of $\phi=0.822$. We want to emphasize that the acceleration of the vibrator was held constant. This means that by changing the frequency $f$, the amplitude of the oscillations was also altered. In Fig.6, we present the transverse spectrum for four different driving frequencies: $f=100,110,120,130$ Hz, with a constant packing fraction of $\phi=0.822$ and acceleration $a=2.5g$.
We observe that the position of the $k$-gap is largely independent of the driving frequency, at least in the range we considered. In contrast, the dispersion of shear waves is sensitive to the driving frequency and exhibits a complex dependence on the frequency $f$. A similar pattern is seen in the dispersion of longitudinal phonons, as illustrated in Fig. 7. Further experimental investigation is necessary to thoroughly analyze how collective motion varies with driving frequency and amplitude.
Although we cannot conduct experiments with frictionless granular particles for direct comparison, we believe that interparticle friction is less significant than the interactions resulting from inelastic collisions between particles. Several pieces of evidence from our previous study of monodisperse particles [49, 50] support this hypothesis.
First, we observed no correlations between individual particles’ translational and rotational degrees of freedom. This suggests that particle rotation due to interparticle friction is largely independent of their translational motions. Second, we successfully explained the flocking behavior observed in our previous monodisperse system [50] by applying the theory of active Brownian particles [58], which completely disregards interparticle friction, as the theory assumes.
Third, we employed vertical driving instead of the more conventional shear driving, reducing interparticle friction’s impact. Lastly, even in our bidisperse systems at the highest packing fraction, the system remains unjammed, indicating that the dominant interactions are due to interparticle collisions rather than contact forces associated with permanent contacts, as seen in jammed solids. Therefore, we anticipate that the results will be comparable to those of frictionless particles.
Figure 6: Transverse dispersion relation under different driving frequency. The transverse dynamical structure factors under different driving frequencies $f=100,110,120,130$ Hz from (a) to (d). The packing fraction is fixed to $\phi=0.822$, and the acceleration is fixed to $a=2.5g$.
Figure 7: Longitudinal dispersion relation under different driving frequency. The longitudinal dynamical structure factor under different driving frequencies $f=100,110,120,130$ Hz from (a) to (d). The packing fraction is fixed to $\phi=0.822$ and the acceleration to $a=2.5g$.
@@ Conclusion
Our study demonstrates that, despite significant differences in the size of composed particles and the absence of a classical thermodynamic description, granular matter on the Brownian vibrator exhibits various similarities with classical liquids in the supercritical region both in the collective and particle-level motions. These alike aspects include pair correlation functions, velocity autocorrelation, mean square displacement on particle levels, vibrational density of states, and dispersion relation of collective excitations. With an increase in the packing fraction $\phi$, distinct phases from gas-like to condensed states (liquid or solid) have been observed on both particle and collective motion levels, suggesting the role of $\phi$ as an effective inverse temperature variable, that is corroborated by the anti-correlation between the average kinetic energy and the packing fraction. However, it is needed to notice that the analogy of temperature is only vague and intuitive and is limited to our specific system. Furthermore, our work experimentally revealed the emergence of the $k$-gap, as predicted by viscoelasticity theory, in granular fluids. This provides a direct link between classical liquid and active granular matter and experimentally confirms the phenomenon of $k$-gap and its theoretical premises.
1 Methods
@@
Unless indicated otherwise, the frequency $f$ is set at 100 Hz, and the maximum acceleration $a$ is 2.5 times the gravitational acceleration $g=9.8m/s^{2}$. The vibration amplitude $A$, defined as $a/(2\pi f)^{2}$, is 0.062 mm, significantly smaller than the particle’s vertical dimension (6 mm). Therefore, we can neglect the vertical displacement of a particle, treating the system as quasi-two-dimensional [49].
@@
These ratios are derived from jamming studies of binary disks, particularly the research conducted by O’hern et al. [76, 77]. Their simulations employed a size ratio of 1:1.4 because a large disk can have up to seven nearest neighbor small disks, while a small disk can have up to five nearest neighbor large disks. This configuration aligns with the theoretical concept of the 7-5 defect roles in the KTHNY theory [78].
In the original O’hern algorithm, a 1:1 number ratio was utilized. However, it was later discovered that a 2:1 number ratio between small and large disks, which corresponds to an area ratio of 1:1, yields even better results. A recent experimental study on the vibrational density of states in jammed 2D granular packing [39] analyzed how the density of states changed when the number ratio varied from a hexagonal crystal (1:0) to the most disordered binary mixture (2:1). The density of states for a system with a 2:1 ratio of small to large disks exhibited several features remarkably similar to those found in molecular glasses [39, 79].
The particles are placed on top of an aluminum plate (60 cm × 60 cm) and confined within a flower-shaped boundary designed to suppress creep particle motions along the boundary [49]. We initiate the experiment with all particles randomly and uniformly placed on the base plate. After applying vibration for two hours, we achieve an initial state of particle configuration. Subsequently, the particle layer undergoes continuous vertical vibration, while a Basler CCD camera (acA2040-180kc) records particle motion at 40 frames/s for at least an hour.
@@
The following fitting functions have been used to analyze the experimentally obtained dynamical structure factor (see additional figures in the Supplementary note 3). The transverse and longitudinal components of the dynamical structure factor are respectively fitted using
@@
The first terms are simply a damped harmonic oscillator with energy $\Omega_{L,T}(k)$ and linewidth $\Gamma_{L,T}(k)$. The second term is a quasi-elastic contribution modelled with a Lorentzian function. The dispersion relations $\Omega_{T,L}(k)$ are shown in Fig.5 as colored bullets and the corresponding linewidths $\Gamma_{T,L}$ by the related errors bars. $S_{\text{loc}}(\omega,k)$ is a term representing possible quasi-localized low-energy modes [81], and it is not relevant for the present discussion. More details on the fitting procedure and the raw data for the dynamical structure factors can be found in the Supplementary note 3.
2 Data availability
The datasets generated and analysed during the current study are available upon reasonable request by contacting the corresponding authors.
3 Acknowledgments
CJ and MB acknowledge the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01). MB acknowledges the support of the sponsorship from the Yangyang Development Fund. ZZ, YC, and JZ acknowledge the support of the NSFC (No. 11974238 and No. 12274291) and the Shanghai Municipal Education Commission Innovation Program under No. 2021-01-07-00-02-E00138. ZZ, YC, and JZ also acknowledge the support from the Shanghai Jiao Tong University Student Innovation Center.
4 Author contributions statement
M. B. and J. Z. conceived the idea of this work and supervised it. Z. Z and Y. C performed the experiments. C. J. performed the data analysis and the theoretical modeling with the help of M. B. All authors contributed to the writing of the manuscript.
5 Competing interests
Jie Zhang is an Editorial Board Member for Communications Physics, but was not involved in the editorial review of, or the decision to publish this article. All the other authors declare no competing interests.
6 Preprints
A preprint of this article is available as, arXiv:2403.08285.
Supplementary notes
Supplementary note 1: Debye normalized density of states
@@ 6 Preprints
$g_{\text{Debye}}(\omega)=\frac{L^{2}}{2\pi\bar{v}^{2}}\,\omega\qquad\text{with}\qquad\frac{2}{\bar{v}^{2}}=\frac{1}{v_{T}^{2}}+\frac{1}{v_{L}^{2}}.$ (S1)
The reduced VDOS $g(\omega)/\omega$ is presented for the two highest values of the packing fraction in Fig.S1.
For the highest packing fraction, $\phi=0.822$, the longitudinal component of the reduced VDOS approaches a constant plateau at low frequency, implying the validity of the Debye model. On the contrary, the transverse component follows Debye’s law only between approximately $10$ Hz to $5$ Hz and then strongly deviates from it below such a frequency. This manifests a residual quasi-elastic diffusive contribution and a finite value of $g_{T}(\omega)$ at zero frequency. By extracting the speeds of longitudinal and transverse sound from the dispersion relation of the collective modes, we have verified that the ratio of the two plateau values in the reduced VDOS (dashed horizontal lines in Fig.S1) is compatible with the theoretical prediction of Debye theory, namely $v_{T}^{2}/v_{L}^{2}$.
Figure S1: The reduced VDOS for packing fraction $\phi=0.811$ (a) and $\phi=0.822$ (b). The blue lines indicate the longitudinal component, while the red and the gray indicate respectively the transverse and total ones. The dashed lines show the theoretical predictions from the Debye model as explained in the text.
The situation becomes more complex by decreasing the packing fraction, as shown in the left panel of Fig.S1 for $\phi=0.811$. There, Debye’s model no longer provides a good description of the experimental data. We can only observe a small plateau in the longitudinal component of the reduced VDOS around $5$ Hz, which is then modified at lower frequencies. In general, by increasing the system’s fluidity or equivalently decreasing its packing fraction, we no longer expect the Debye model to be a good representation of the collective dynamics.
Supplementary note 2: Static structure factor
We have obtained the experimental structure factor $S(k)$ for four different packing fractions to provide further experimental evidence for a connection between the dynamical Frenkel line and structural changes. The results are shown in Fig. S2. From there, a gradual disappearance of the medium-range structural order is evident by decreasing $\phi$ from the maximal value $\phi=0.822$ to a low value $\phi=0.618$. In particular, by reducing $\phi$, we notice that (I) the intensity of the first peak around $k\approx d_{s}^{-1}$ diminishes, (II) the depth of the valley around $k\approx 1.5d_{s}^{-1}$ becomes smaller and (III) the higher order peaks gradually disappear. These results imply a gradual loss of medium-range order from the dense liquid phase (large $\phi$) to the gas-like dilute phase (small $\phi$).
Figure S2: The experimental static structure factor for different values of the packing fraction $\phi$.
Supplementary note 3: Dynamical structure factor
We provide more details about the analysis of the experimental dynamical structure factor. The longitudinal dynamical structure factor is fitted using a single damped harmonic oscillator,
$S_{L}(\omega,k)\propto\dfrac{\omega^{2}}{(\omega^{2}-\Omega_{L}^{2})^{2}+\omega^{2}\Gamma_{L}^{2}},$ (S2)
@@
$S_{T}(\omega,k)\propto\dfrac{\omega^{2}}{(\omega^{2}-\Omega_{T}^{2})^{2}+\omega^{2}\Gamma_{T}^{2}}+\dfrac{1}{\pi}\dfrac{\gamma}{\omega^{2}+\gamma^{2}}+S_{loc}(\omega,k),$ (S3)
@@
$S_{loc}(\omega,k)=\dfrac{1}{\omega\sigma\sqrt{2\pi}}\mathrm{exp}\left(-\dfrac{(\mathrm{ln}\omega-\mu)^{2}}{2\sigma^{2}}\right).$ (S4)
@@
Supplementary note 4: Displacement field power spectrum
@@
We have performed the Fourier transform on the displacement field to observe its evolution over time. Fig. S6 presents the results for different packing fractions. For low packing fractions (a) and (b), all Fourier transform components concentrate strongly and stably in the low wave vector region after a long observation time (800s). This indicates that the particles do not return to their initial position. However, the features differ for high packing fractions, as shown in panels (c) and (d). First, the Fourier components are unstable, especially the zigzag structures at $k\approx 0$. This suggests that the particles may still return to their initial position. Secondly, the Fourier transform components concentrate at a wave vector smaller than $k$-gap. This correlation can be interpreted as follows: the modes with $k>k_{g}$ are solid-like, meaning that the particles only vibrate around their equilibrium position, and therefore these modes do not contribute to the displacement at a late time. On the other hand, the modes with $k<k_{g}$ are gas- or liquid-like, meaning that the particles can diffuse away from their position and never return. Therefore, only these modes contribute to the long-time displacement.
Supplementary note 5: Lifetime of local atomic connectivity
@@
We calculate the lifetime of local atomic connectivity, $\tau_{LC}$, which represents the average time it takes for atoms to lose or gain one nearest neighbor, changing the local atomic connectivity through a local configurational excitation. The Voronoi cell of the particle at time $t=t_{0}$ defines its nearest neighbors. The schematic diagram of Fig.S7 indicates that any two particles that share an edge of a Voronoi cell are considered neighbors, as explained in detail in [82]. In doing so, the atomic bonds of a particle with its neighbors can be defined. After that, we can detect the formation or breaking of atomic bonds for $t>t_{0}$ by identifying the moments when the number of neighbors changes by a unity. For each atom, $t_{LC}$ is the time to change one neighbor such that $N_{c}(t_{0}+t_{LC})=N_{c}(t_{0})\pm 1$. The lifetime of local atomic connectivity is then defined as the average of $t_{LC}$, which is denoted as $\tau_{LC}=\langle t_{LC}\rangle$. The distribution of $t_{LC}$ for different packing fractions is shown in Fig.S8. $\tau_{LC}$ values for the $\phi$ values 0.618, 0.773, 0.811, and 0.822 are 0.308, 0.609, 0.507, and 0.300s, respectively.
@@
Figure S3: Transverse (red,a) and longitudinal (blue,b) dynamical structure factor at different wave vector for $\phi=0.822$. The thinner lines in transverse panel show the three terms in the fitting function.
@@
Figure S4: Transverse (red,a) and longitudinal (blue,b) dynamical structure factor at different wave vector for $\phi=0.811$. The thinner lines in transverse panel show the three terms in the fitting function.
@@
Figure S5: The fitting parameters $\Omega$ (circle) and $\Gamma$ (square) for longitudinal (blue) and transverse (red) dynamical structure factor and for packing fraction $\phi=0.811$ (a) and $\phi=0.822$ (b).
@@
Figure S6: Fourier transform on displacement field, $\boldsymbol{D}(\boldsymbol{k})=\mathcal{F}[\boldsymbol{D}(\boldsymbol{x})]$, for packing fraction $\phi=0.618$, $\phi=0.773$, $\phi=0.811$ and $\phi=0.822$ from (a) to (d) respectively. Since our system have rotational symmetry, only $x$ component $D_{x}(\boldsymbol{x})$ are plot. The darker color indicate stronger Fourier transform signal. The red lines in panels (c) and (d) show the position of $k$-gap.
Figure S7: Schematic diagram for an atom to change one neighbor.
Figure S8: Distribution of $t_{LC}$ for packing fraction $\phi=0.618$, $\phi=0.773$, $\phi=0.811$ and $\phi=0.822$ from (a) to (d). $\tau_{LC}=0.308,0.609,0.507,0.300s$ for each packing fraction respectively.
Supplementary note 6: Dynamics of a single active particle
In Fig. S9, we show two typical trajectories for one small particle and one large particle under an observation time of $600$ s (from blue to red color).
The MSD and RMSD for small and large particles are, respectively, shown in panels (left) and (right) of Fig.S10. The left panel demonstrates that the MSD curves of both small and large particles are almost identical, exhibiting both short-time quasi-ballistic and long-time diffusive motions. The quasiballistic motion observed at short time scales corresponds to a persistent time of approximately $\tau_{p}\approx 0.25$ seconds for both small and large particles. The diffusion process begins for $\tau\geq\tau_{p}$ on the MSD curves, as illustrated in Fig.S10(left). From the long-time diffusive motions, we derive their translational diffusion constants $D\approx 0.45d_{s}^{2}$/s for both small and large particles.
The right panel illustrates that the RMSD curves for both small and large particles are nearly indistinguishable, indicating weakly super-diffusive motions. This is evidenced by a slope close to 1.4, as shown by the triangle on the log-log plot. The observed super-diffusive behavior arises from the technical challenges associated with manufacturing perfectly rotationally symmetric particles. This behavior can be quantified using the fitting function $\text{RMSD}=D_{r,{\ast}}t^{\gamma}$, with $\ast$ representing ‘l’ and ‘s’, yielding the constants $D_{r,l}=9.12$, $D_{r,s}=9.33$, $\gamma_{l}=1.43$, and $\gamma_{s}=1.35$, where the subscripts ‘l’ and ‘s’ refer to ‘large’ and ‘small’ particles, respectively.
The single-particle dynamics clarify the active nature of our experimental setup.
Figure S9: Typical trajectories of small and large particles. The color indicates the time evolution within an observation window of $600$ seconds.
Figure S10: The mean square displacement MSD (left) and rotational mean square displacement RMSD (right) based on single particle experiment for both small and large particles.